OpenStudy (anonymous):

Will givve a medals Provide the solution to the equation 25 = square root of 5 (2x + 6).Part 2: Using complete sentences, explain the strategy used in solving this equation the way you did. (3 points)

OpenStudy (ankit042):

Sqaure both the sides 25*25 = 5(2x+6) 125 = 2x+6 119/2 =x

OpenStudy (anonymous):

@ankit042 can u explain how u got thaat?